# ADEPT: Adaptive Diffusion Environment for Policy Transfer Sim-to-Real

### v1

@@ V-A Algorithmic Performance Evaluation
ADEPT generates realistic environments. The higher success rate of ADEPT than APG on the real-world replicated environments show the generation quality of ADEPT empowers robot navigation policy learning, compared to AAPG and N-AEPT in our recorded data. In the following sim-to-sim and sim-to-real experiments, we will demonstrate the smooth motion trained through ADEPT compared to the well-performing but unnatural policy from procedural generations.
ADEPT evolves environment difficulty. The RL return curve reflects the stable performance of ADEPT on evaluation environments, attributed to the evolving difficulty of training environments generated by ADEPT. As the policy encounters progressively harder environments, its performance initially dips but gradually stabilizes and converges. Although N-AEPTPT enjoys a large training dataset, it can hardly outperform ADEPT due to the lack of difficulty controllability. PG’s sharp curve changes due to overly difficult environments also show the importance of ADEPT.
ADEPT evolves environment diversity. ADEPT gains advantages over a fixed dataset such as N-AEPTPT because ADEPT can easily generate thousands of environments within tens of epochs. Procedural Generation is limited by the parametric range and lacks efficient environment parameter control.
@@ V-C Simulation Experiment
We simulate wheeled robots, ClearPath Jackal, in ROS Gazebo on 30 diverse environments (E-30), equipped with a RealSense D435 camera ($30$ Hz). We add Gaussian noises to the ground-truth robot state ($200$ Hz), depth measurement, and vehicle control to introduce uncertainty whose parameters reflect the hardest curriculum during simulation training. The ROS message filter synchronized the odometry with depth measurement. $1000$ start and goal pairs are sampled for each environment. We do not include ablations other than N-AEPT because of poor algorithmic performance. As results shown in Table I, our method outperforms the baselines. While all methods show improved performance due to the Husky’s better navigability on uneven terrains, our method consistently outperformed baseline methods. The depth measurement noise poses a substantial challenge in accurately modeling obstacles and complex environments. Falco and MPPI often cause the robot to get stuck or topple over, and TERP often predicts erratic waypoints that either violate safety on elevation map or are overly conservative. Learning-based TERP and POVN lack generalizability, with their performance varying across different environments. This issue is mirrored in N-AEPT and APG, highlighting the success of adaptive curriculum and realistic environment generation properties of ADEPT.

### v2

@@ V-A Algorithmic Performance Evaluation
ADEPT generates realistic environments. The higher success rate of ADEPT than APG on the real-world replicated environments show the generation quality of ADEPT empowers robot navigation policy learning, compared to APG and N-AEPT. In the following sim-to-sim and sim-to-real experiments, we will demonstrate the smooth motion trained through ADEPT compared to the well-performing but unnatural policy from procedural generations.
ADEPT evolves environment difficulty. The RL return curve reflects the stable performance of ADEPT on evaluation environments, attributed to the evolving difficulty of training environments generated by ADEPT. As the policy encounters progressively harder environments, its performance initially dips but gradually stabilizes and converges. Although N-AEPT enjoys a large training dataset, it can hardly outperform ADEPT due to the lack of difficulty controllability.
ADEPT evolves environment diversity. ADEPT gains advantages over a fixed dataset such as N-AEPT because ADEPT can easily generate thousands of environments within tens of epochs. PG is limited by the parametric range and lacks efficient environment parameter control.
In summary, ADEPT excels at adapting environment difficulty and diversity based on evolving policy performance.
@@ V-C Simulation Experiment
We simulate wheeled robot, ClearPath Jackal, in ROS Gazebo on 30 diverse environments (E-30), equipped with a RealSense D435 camera ($30$ Hz). We add Gaussian noises to the ground-truth robot state ($200$ Hz), depth measurement, and vehicle control to introduce uncertainty whose parameters reflect the hardest curriculum during simulation training. The ROS message filter synchronized the odometry with depth measurement. $1000$ start and goal pairs are sampled for each environment. We do not include ablations other than N-AEPT because of poor algorithmic performance. As results shown in Table I, our method outperforms the baselines. While all methods show improved performance due to the Husky’s better navigability on uneven terrains, our method consistently outperformed baseline methods. The depth measurement noise poses a substantial challenge in accurately modeling obstacles and complex environments. Falco and MPPI often cause the robot to get stuck or topple over, and TERP often predicts erratic waypoints that either violate safety on elevation map or are overly conservative. Learning-based TERP and POVN lack generalizability, with their performance varying across different environments. This issue is mirrored in N-AEPT and APG, highlighting the success of adaptive curriculum and realistic environment generation properties of ADEPT.